OpenStudy (anonymous):

Write each fraction as a decimal.3 - 4 show me how reamember i'm a 5th grader.

OpenStudy (jagatuba):

You should post this question to the Math group and your questions is a bit hard to read. Are you wanting to know how to write 3/4 or \[\frac{3}{4}\]as I think you were trying to do? One way to look at fractions is to see them as division problems. \[\frac{3}{4}=3\div4=0.25\]Anytime that you have a fraction all you have to do is divide the top number by the bottom number and you will get your decimal equivalent. Be careful with some fractions though because they repeat forever. \[\frac{1}{3}=1\div3=0.33333333333333.......\] If you have not already learned about repeating decimals, you will soon.